OpenStudy (anonymous):

Find the exact value of cot-1(1).

OpenStudy (anonymous):

Please help me!

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

\[arccot(1)=\theta\]\[\cot \theta=1\] The inverse function produces an angle, so we can write that as theta. From there we can rewrite it as cotangent of theta. And once again we have to think about our special angles D: This one is a little tricky.

OpenStudy (anonymous):

45 degrees = pi/4

OpenStudy (anonymous):

@ballerinaxbb

zepdrix (zepdrix):

...

OpenStudy (anonymous):

Hi sorry

OpenStudy (anonymous):

I'm back

OpenStudy (anonymous):

your question is done

OpenStudy (precal):

your angle is pi/4 or 45 degrees

OpenStudy (anonymous):

Thanks to both of you! @zepdrix please do me a favor and give @vladislavchazov a medal too :)

OpenStudy (anonymous):

Thanks @precal :)

OpenStudy (anonymous):

your welcome @ballerinaxbb

OpenStudy (precal):

yw